OpenStudy (anonymous):

Solve 2x + 5y = –13 3x – 4y = –8 (4, 1) (–4, 1) (4, –1) (–4, –1)

OpenStudy (anonymous):

@precal

OpenStudy (precal):

just plug in each value until both are true recall (x,y) example (4,1) means x=4 and y=1 just sub into both equations, if it is true then that is the solution

OpenStudy (precal):

Systems of equations are 2 lines that might cross. That cross point is a solution. In this case, you have 2 equations (lines) that are crossing at one point. You are being asked to find the point of intersection.

OpenStudy (anonymous):

@precal I Got B.

OpenStudy (anonymous):

set them equal to each other

OpenStudy (anonymous):

@callie2240 I Got B. Is That Correct?

OpenStudy (anonymous):

I have to solve it first to know for sure lol

OpenStudy (anonymous):

Multiply equation 1 by 3 and you get 6x + 15y = -39 Multiply equation 2 by 2 and you get 6x - 8y = -16 Subtract the second from the first .... 23y = -23 Therefore y = -1 Substitute y=-1 into 2x + 5y = -13 and we get 2x - 5 = -13 so 2x = -8 so x = -4 So the answer is (-4, -1) so no, the answer is d understand?